OpenStudy (amber15):

What is the perimeter of a computer graphic woth sides measuring 2 sqaureroot 27 . 6 sqaure root 48 , 5 squareroot 75 and 6 sqaure root 20 centimeters . the perimeter is ---- cemtimeters

OpenStudy (anonymous):

\[2 \sqrt{27}+6 \sqrt{48}+5 \sqrt{75}+6 \sqrt{20}=55 \sqrt{3}+12 \sqrt{5}=122.096 \]

OpenStudy (amber15):

okay

OpenStudy (anonymous):

\[6 \sqrt{3}+24 \sqrt{3}+25 \sqrt{3}+12 \sqrt{5}=55 \sqrt{3}+12 \sqrt{5} \]

OpenStudy (amber15):

Thank you